For each structure below, determine the number of
Answer #1
Concepts and Reason

The problem is based on the concept of aromaticity of organic compounds and Huckel’s rule. An aromatic compound represents a substance, which comprises of one or more than one rings with alternating single and double bonds. Aromatic compounds follow the Huckel’s rule, which describes that a compound to be aromatic must contain (4n+2)\left( {4n + 2} \right) number of π\;\pi electrons here, n is any integer.

Fundamentals

An organic compound is said to be aromatic if it satisfies the following conditions:

1. It must have cyclic structure.

2. In the cyclic structure it must have conjugated double bonds

3. It must follow Huckel’s rule.

A nonaromatic compound is represented by the non-planar or non-cyclic structure. Lastly, an antiaromatic compound represents a cyclic substance, which has an π\;\pi electron system containing 4n4n electrons. Where, n is any integer.

(a)

In this case, the structure contains 3π3\pi bonds. Thus, the number of π\;\pi electrons is 66

(a)

The ring structure is an aromatic compound.

(b)

In this case, the structure contains 2π\;2\pi bonds. Thus, the number of π\;\pi electrons is 44

(b)

The structure is an antiaromatic compound.

Ans: Part a

The number of π\;\pi electrons in the given structure is 66 and it indicates an aromatic structure.